# Physics-Aware, Shannon-Optimal Compression via Arithmetic Coding for Distributional Fidelity

### v2

@@ Abstract
Assessing whether two datasets are distributionally consistent has become a central theme in modern scientific analysis, particularly as generative artificial intelligence is increasingly used to produce synthetic datasets whose fidelity must be rigorously validated against the original data on which they are trained, a task made more challenging by the continued growth in data volume and problem dimensionality. In this work, we propose the use of arithmetic coding to provide a lossless and invertible compression of datasets under a physics-informed probabilistic representation. Datasets that share the same underlying physical correlations admit comparable optimal descriptions, while discrepancies in those correlations—arising from miscalibration, mismodeling, or bias—manifest as an irreducible excess in code length. This excess codelength defines an operational fidelity metric, quantified directly in bits through differences in achievable compression length relative to a physics-inspired reference distribution. We demonstrate that this metric is global, interpretable, additive across components, and asymptotically optimal in the Shannon sense. Moreover, we show that differences in codelength correspond to differences in expected negative log-likelihood evaluated under the same physics-informed reference model. As a byproduct, we also demonstrate that our compression approach achieves a higher compression ratio than traditional general-purpose algorithms such as gzip. Our results establish lossless, physics-aware compression based on arithmetic coding not as an end in itself, but as a measurement instrument for testing the fidelity between datasets.
@@ III.1 Arithmetic Coding and Shannon Optimality
Consider data drawn from an unknown distribution $p(x)$ that are encoded using a (possibly mismatched) model $q(x)$. Taking expectations under $p$, the average achieved length converges to the cross-entropy,
$\mathbb{E}_{x\sim p}\!\big[\ell_{q}(x)\big]\;\simeq\;\mathbb{E}_{x\sim p}\!\big[-\log_{2}q(x)\big]\;=\;H(p)+D_{\mathrm{KL}}(p\|q),$ (5)
Taking expectations under $p$, the average achieved length converges to the cross-entropy between $p$ and $q$,
@@ III.3 Lossless Compression as a Likelihood-Scored Fidelity Diagnostic
Arithmetic coding provides an operational realization of negative log-likelihood under a fixed reference model. Let $q_{A}$ denote the probability model (CDF tables) learned from a reference dataset AA and then held fixed. For a dataset $D=\{x_{i}\}_{i=1}^{N}$ encoded with this codec, the total codelength satisfies
@@
Here $\hat{p}_{D}$ denotes the empirical distribution associated with dataset $D$, while the unknown population distribution is denoted by pp. Given an independent baseline sample $B$ from the same population as AA and a perturbed (or synthetic) sample $C_{\varepsilon}$, we define the excess codelength
@@ IV.1 Invertibility
To verify the lossless and invertible nature of the proposed arithmetic coding framework, we perform a direct comparison between detector-level observables computed from the original data and from the data obtained after a full compression–decompression cycle. The dataset $R$ is split into disjoint training (AA) and validation ($B$) subsets, with models trained on AA and evaluated on $B$. Figure 1 reports representative results for dataset $B$.
@@ IV.3 Comparison between Shannon Entropy and Average Sequence Length
Tables 2 and 3 report the empirical entropy $H(q)$, the cross-entropy $H(p,q)$, and the achieved average code length $L$ (all in bits per event) for the unconditional and conditional codecs, respectively. Results are shown separately for the calorimeter hits and for the particle-kinematics payload, as well as for the total event representation, and for both the training dataset AA and the validation $B$. Several features are worth highlighting.
@@
Second, the cross-entropy evaluated on the validation dataset $B$ is systematically larger than that obtained on the training dataset AA. This increase is consistent with the expected additional “surprise” incurred when encoding previously unseen data drawn from the same underlying process, reflecting finite-sample estimation effects rather than a true distributional difference, since the reference model is trained on AA rather than on $B$ [18].
@@ IV.4 Information and Bit-Budget Decomposition
Figure 2 shows the achieved bit budget for the unconditional arithmetic coding model trained on split A, decomposed by calorimeter layer and stereo view. For each layer–view (LV), the total contribution is further separated into occupancy, strip index, and ADC amplitude terms.
@@
Table 4 reports the corresponding numerical values. The dominant contribution arises from ADC amplitudes, followed by strip indices, while occupancy carries a smaller but non-negligible fraction of the total information. The PCAL layers contribute the largest share of the bit budget, reflecting both higher hit multiplicities and a broader ADC amplitude range, while ECIN and ECOUT contribute progressively less. The final column reports the mean hit multiplicity per layer–view, illustrating the close correspondence between detector activity and information content. Summing over all nine layer–views yields a detector-level contribution of approximately $779.55$ bits per event, which can be compared to Table 2. The particle kinematic information, encoded separately using a generic byte-level model, contributes an additional $82.43$ bits per event. The resulting total achieved codelength of $\approx 862$ bits per event defines the unconditional reference baseline. Any possible reduction in codelength achieved by conditional or enriched probabilistic models can therefore be interpreted directly as recovered physical correlations, rather than changes in the coding algorithm.
@@ IV.5 Fidelity Studies
Let $q_{A}$ denote the probabilistic model (CDFs) learned from the reference sample AA and then held fixed. Following Eqs. (11) and (12), for a dataset $D=\{x_{i}\}_{i=1}^{N}$ the arithmetic codec induces an average per-event codelength (in bits per event), $L_{A}(D)$, up to negligible termination overhead (Sec. IV.3). We define the excess codelength
@@
evaluated under the same fixed reference codec trained on AA. Let $\mu_{C}(\varepsilon)$ and $\mu_{B}$ denote the sample means of $\{L_{A}(C_{\varepsilon,k})\}$ and $\{L_{A}(B_{k})\}$, respectively, with sample standard deviations $s_{C}(\varepsilon)$ and $s_{B}$. We define the mean excess codelength

### v3

@@ Abstract
Assessing whether two datasets are distributionally consistent is central to modern scientific analysis, particularly as generative artificial intelligence produces synthetic data whose fidelity must be validated against real observations in increasingly high-dimensional settings. Existing approaches are typically relative: they determine whether one dataset is more consistent with a reference than another, but do not provide a physically grounded absolute standard for fidelity. We propose an information-theoretic approach in which lossless compression via arithmetic coding provides an operational measure of dataset fidelity under a physics-informed probabilistic representation. Datasets sharing the same underlying physical correlations admit comparable optimal descriptions, while discrepancies-arising from miscalibration, mismodeling, or bias-manifest as an irreducible excess in codelength relative to the Shannon-optimal limit defined by the physics itself. This excess codelength defines an absolute fidelity metric, quantified directly in bits. Unlike conventional measures, which lack an intrinsic scale, zero excess provides a well-defined and physically meaningful target corresponding to consistency with the underlying distribution. We show that this metric is global, interpretable, additive across components, and asymptotically optimal, with differences in codelength corresponding to differences in expected negative log-likelihood under a common reference model. As a byproduct, our approach achieves improved compression relative to standard general-purpose algorithms such as gzip. These results establish arithmetic coding not merely as a compression tool, but as a measurement instrument for absolute, physics-grounded assessment of distributional fidelity.
@@ III.1 Arithmetic Coding and Shannon Optimality
Consider data drawn from an unknown distribution $p(x)$ that are encoded using a (possibly mismatched) model $q(x)$. Taking expectations under $p$, the average achieved length converges to the cross-entropy between $p$ and $q$,
@@ III.3 Lossless Compression as a Likelihood-Scored Fidelity Diagnostic
Arithmetic coding provides an operational realization of negative log-likelihood under a fixed reference model. Let $q_{A}$ denote the probability model (CDF tables) learned from a reference dataset $A$ and then held fixed. For a dataset $D=\{x_{i}\}_{i=1}^{N}$ encoded with this codec, the total codelength satisfies
@@
Here $\hat{p}_{D}$ denotes the empirical distribution associated with dataset $D$, while the unknown population distribution is denoted by $p$. Given an independent baseline sample $B$ from the same population as $A$ and a perturbed (or synthetic) sample $C_{\varepsilon}$, we define the excess codelength
@@ IV.1 Invertibility
To verify the lossless and invertible nature of the proposed arithmetic coding framework, we perform a direct comparison between detector-level observables computed from the original data and from the data obtained after a full compression–decompression cycle. The dataset $R$ is split into disjoint training ($A$) and validation ($B$) subsets, with models trained on $A$ and evaluated on $B$. Figure 1 reports representative results for dataset $B$.
@@ IV.3 Comparison between Shannon Entropy and Average Sequence Length
Tables 2 and 3 report the empirical entropy $H(q)$, the cross-entropy $H(p,q)$, and the achieved average code length $L$ (all in bits per event) for the unconditional and conditional codecs, respectively. Results are shown separately for the calorimeter hits and for the particle-kinematics payload, as well as for the total event representation, and for both the training dataset $A$ and the validation $B$. Several features are worth highlighting.
@@
Second, the cross-entropy evaluated on the validation dataset $B$ is systematically larger than that obtained on the training dataset $A$. This increase is consistent with the expected additional “surprise” incurred when encoding previously unseen data drawn from the same underlying process, reflecting finite-sample estimation effects rather than a true distributional difference, since the reference model is trained on $A$ rather than on $B$ [18].
@@ IV.4 Information and Bit-Budget Decomposition
Figure 2 shows the achieved bit budget for the unconditional arithmetic coding model trained on split A, decomposed by calorimeter layer and stereo view. For each layer–view (LV), the total contribution is further separated into occupancy, strip index, and ADC amplitude terms. Table 4 reports the corresponding numerical values. The dominant contribution arises from ADC amplitudes, followed by strip indices, while occupancy carries a smaller but non-negligible fraction of the total information. The PCAL layers contribute the largest share of the bit budget, reflecting both higher hit multiplicities and a broader ADC amplitude range, while ECIN and ECOUT contribute progressively less. The final column reports the mean hit multiplicity per layer–view, illustrating the close correspondence between detector activity and information content. Summing over all nine layer–views yields a detector-level contribution of approximately $779.55$ bits per event, which can be compared to Table 2. The particle kinematic information, encoded separately using a generic byte-level model, contributes an additional $82.43$ bits per event. The resulting total achieved codelength of $\approx 862$ bits per event defines the unconditional reference baseline. Any possible reduction in codelength achieved by conditional or enriched probabilistic models can therefore be interpreted directly as recovered physical correlations, rather than changes in the coding algorithm.
@@ IV.5 Fidelity Studies
Let $q_{A}$ denote the probabilistic model (CDFs) learned from the reference sample $A$ and then held fixed. Following Eqs. (10) and (11), for a dataset $D=\{x_{i}\}_{i=1}^{N}$ the arithmetic codec induces an average per-event codelength (in bits per event), $L_{A}(D)$, up to negligible termination overhead (Sec. IV.3). We define the excess codelength
@@
evaluated under the same fixed reference codec trained on $A$. Let $\mu_{C}(\varepsilon)$ and $\mu_{B}$ denote the sample means of $\{L_{A}(C_{\varepsilon,k})\}$ and $\{L_{A}(B_{k})\}$, respectively, with sample standard deviations $s_{C}(\varepsilon)$ and $s_{B}$. We define the mean excess codelength